RMEX has a market capitalization of $100 million and $40 million in debt. RMEX intends to maintain the same capital structure in the future. The corporate tax rate is 33%. If RMEX's anticipated operating cash flow (OCEF) for next year is $7 million and its anticipated growth is 3% per year to infinity. What is RMEX's after-tax weighted average cost of capital (WACC)?
Answer #1

If company plans to maintain same capital structure in future

Value of firm = FCF/(WACC -g)

Value of firm = Value of equity + value of debt = 100 + 40 = 140

140 = 7/(WACC - 0.03)

(WACC - 0.03) = 7/140 = 0.05

WACC = 0.05 + 0.03 =0.08 = 8%

Answer = 8%